# and thermofield double states as quantum many-body scars

### v3

@@ Acknowledgements.
Supplementary material
@@
$|O\rangle\rangle_{0}=O\otimes\mathbb{1}\;|\mathbb{1}\rangle\rangle_{0}=\mathbb{1}\otimes O^{\mathsf{T}}\;|\mathbb{1}\rangle\rangle_{0}.$ (S3)
@@
$|O\rangle\rangle_{\beta}=\sum_{m,n}O_{mn}e^{-\frac{\beta}{4}H_{0}}|m\rangle\otimes e^{-\frac{\beta}{4}H_{0}^{T}}|n\rangle=e^{-\frac{\beta}{4}H_{0}}\otimes e^{-\frac{\beta}{4}H_{0}^{\mathsf{T}}}\;|O\rangle\rangle_{0}=e^{-\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{\mathsf{T}})}|O\rangle\rangle_{0}.$ (S4)
@@
$|O\rangle\rangle_{\beta}=(e^{-\frac{\beta}{4}H_{0}}\otimes e^{-\frac{\beta}{4}H_{0}^{\mathsf{T}}})(O\otimes\mathbb{1})|\mathbb{1}\rangle\rangle_{0}$ $=e^{-\frac{\beta}{4}H_{0}}O\otimes e^{-\frac{\beta}{4}H_{0}^{\mathsf{T}}}\;|\mathbb{1}\rangle\rangle_{0}$
$=e^{-\frac{\beta}{4}H_{0}}Oe^{-\frac{\beta}{4}H_{0}}\otimes\mathbb{1}\;|\mathbb{1}\rangle\rangle_{0}=\mathbb{1}\otimes e^{-\frac{\beta}{4}H_{0}^{\mathsf{T}}}O^{\mathsf{T}}e^{-\frac{\beta}{4}H_{0}^{\mathsf{T}}}\;|\mathbb{1}\rangle\rangle_{0}.$ (S5)
@@
$|\mathbb{1}\rangle\rangle_{\beta}=e^{-\frac{\beta}{2}H_{0}}\otimes\mathbb{1}\;|\mathbb{1}\rangle\rangle_{0}=\mathbb{1}\otimes e^{-\frac{\beta}{2}H_{0}^{\mathsf{T}}}\;|\mathbb{1}\rangle\rangle_{0},$ (S6)
@@
$|O\rangle\rangle_{\beta}=e^{-\frac{\beta}{4}H_{0}}Oe^{\frac{\beta}{4}H_{0}}\otimes\mathbb{1}\;|\mathbb{1}\rangle\rangle_{\beta}=\mathbb{1}\otimes e^{-\frac{\beta}{4}H_{0}^{\mathsf{T}}}O^{\mathsf{T}}e^{\frac{\beta}{4}H_{0}^{\mathsf{T}}}\;|\mathbb{1}\rangle\rangle_{\beta}.$ (S7)
@@
$|AOB\rangle\rangle_{\beta}$ $=e^{-\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{\mathsf{T}})}|AOB\rangle\rangle_{0}$
$=e^{-\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{\mathsf{T}})}A\otimes B^{T}|O\rangle\rangle_{0}=e^{-\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{\mathsf{T}})}A\otimes B^{T}e^{\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{\mathsf{T}})}|O\rangle\rangle_{\beta}.$
@@
$\mathcal{H}_{\beta}=e^{-\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{\mathsf{T}})}\mathcal{H}e^{\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{\mathsf{T}})},$ (S9)
@@
$\mathcal{H}$ $=-H\otimes\mathbb{1}+\mathbb{1}\otimes H^{\mathsf{T}}$
$+\frac{g}{2}\sum_{j}\biggl{(}e^{\frac{\beta}{2}}\left(L_{j}^{\dagger}\otimes L_{j}^{\mathsf{T}}-\frac{1}{2}L_{j}^{\dagger}L_{j}\otimes\mathbb{1}-\frac{1}{2}\mathbb{1}\otimes L_{j}^{\mathsf{T}}L_{j}^{*}\right)+e^{-\frac{\beta}{2}}\left(L_{j}\otimes L_{j}^{*}-\frac{1}{2}L_{j}L_{j}^{\dagger}\otimes\mathbb{1}-\frac{1}{2}\mathbb{1}\otimes L_{j}^{*}L_{j}^{\mathsf{T}}\right)\biggr{)}.$

### v4

@@ Acknowledgements.
Supplementary material for Letter
“Duality between open systems and closed bilayer systems,
and thermofield double states as quantum many-body scars”
by Alexander Teretenkov and Oleg Lychkovskiy
@@
$|O\rangle\rangle_{0}=O\otimes\mathbb{1}\;|\mathbb{1}\rangle\rangle_{0}=\mathbb{1}\otimes O^{T}\;|\mathbb{1}\rangle\rangle_{0}.$ (S3)
@@
$|O\rangle\rangle_{\beta}=\sum_{m,n}O_{mn}e^{-\frac{\beta}{4}H_{0}}|m\rangle\otimes e^{-\frac{\beta}{4}H_{0}^{T}}|n\rangle=e^{-\frac{\beta}{4}H_{0}}\otimes e^{-\frac{\beta}{4}H_{0}^{T}}\;|O\rangle\rangle_{0}=e^{-\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{T})}|O\rangle\rangle_{0}.$ (S4)
@@
$|O\rangle\rangle_{\beta}=(e^{-\frac{\beta}{4}H_{0}}\otimes e^{-\frac{\beta}{4}H_{0}^{T}})(O\otimes\mathbb{1})|\mathbb{1}\rangle\rangle_{0}$ $=e^{-\frac{\beta}{4}H_{0}}O\otimes e^{-\frac{\beta}{4}H_{0}^{T}}\;|\mathbb{1}\rangle\rangle_{0}$
$=e^{-\frac{\beta}{4}H_{0}}Oe^{-\frac{\beta}{4}H_{0}}\otimes\mathbb{1}\;|\mathbb{1}\rangle\rangle_{0}=\mathbb{1}\otimes e^{-\frac{\beta}{4}H_{0}^{T}}O^{T}e^{-\frac{\beta}{4}H_{0}^{T}}\;|\mathbb{1}\rangle\rangle_{0}.$ (S5)
@@
$|\mathbb{1}\rangle\rangle_{\beta}=e^{-\frac{\beta}{2}H_{0}}\otimes\mathbb{1}\;|\mathbb{1}\rangle\rangle_{0}=\mathbb{1}\otimes e^{-\frac{\beta}{2}H_{0}^{T}}\;|\mathbb{1}\rangle\rangle_{0},$ (S6)
@@
$|O\rangle\rangle_{\beta}=e^{-\frac{\beta}{4}H_{0}}Oe^{\frac{\beta}{4}H_{0}}\otimes\mathbb{1}\;|\mathbb{1}\rangle\rangle_{\beta}=\mathbb{1}\otimes e^{-\frac{\beta}{4}H_{0}^{T}}O^{T}e^{\frac{\beta}{4}H_{0}^{T}}\;|\mathbb{1}\rangle\rangle_{\beta}.$ (S7)
@@
$|AOB\rangle\rangle_{\beta}$ $=e^{-\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{T})}|AOB\rangle\rangle_{0}$
$=e^{-\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{T})}A\otimes B^{T}|O\rangle\rangle_{0}=e^{-\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{T})}A\otimes B^{T}e^{\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{T})}|O\rangle\rangle_{\beta}.$
@@
$\mathcal{H}_{\beta}=e^{-\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{T})}\mathcal{H}e^{\frac{\beta}{4}(H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}^{T})},$ (S9)
@@
$\mathcal{H}$ $=-H\otimes\mathbb{1}+\mathbb{1}\otimes H^{T}$
$+\frac{g}{2}\sum_{j}\biggl(e^{\frac{\beta}{2}}\left(L_{j}^{\dagger}\otimes L_{j}^{T}-\frac{1}{2}L_{j}^{\dagger}L_{j}\otimes\mathbb{1}-\frac{1}{2}\mathbb{1}\otimes L_{j}^{T}L_{j}^{*}\right)+e^{-\frac{\beta}{2}}\left(L_{j}\otimes L_{j}^{*}-\frac{1}{2}L_{j}L_{j}^{\dagger}\otimes\mathbb{1}-\frac{1}{2}\mathbb{1}\otimes L_{j}^{*}L_{j}^{T}\right)\biggr).$
@@ C.2 Generalized Hubbard model
Appendix F S.6 First few states in the tower of scars
Here we provide explicit expressions for two scars of the model (14) dual to eigen-HIOMS (Duality between open systems and closed bilayer systems, and thermofield double states as quantum many-body scars ). According to eq. (20), these (unnormalized) scar eigenstates and eigenenergies read
$|Q_{1}\rangle\rangle_{\beta}$ $=|S^{z}\rangle\rangle_{\beta}+\frac{N}{2}\tanh\frac{\beta}{2}\,|{\mathbb{1}}\rangle\rangle_{\beta},$ $E_{1}$ $=-g\cosh\frac{\beta}{2},$
$|Q_{2}\rangle\rangle_{\beta}$ $=|\left(S^{z}\right)^{2}\rangle\rangle_{\beta}+(N-1)\tanh\frac{\beta}{2}|S^{z}\rangle\rangle_{\beta}+\frac{N}{4}\left((N-1)\tanh^{2}\frac{\beta}{2}-1\right)\,|{\mathbb{1}}\rangle\rangle_{\beta},$ $E_{2}$ $=-2\,g\cosh\frac{\beta}{2},$ (S33)
with $|{\mathbb{1}}\rangle\rangle_{\beta}$ given by eq. (8) and
$|S^{z}\rangle\rangle_{\beta}$ $=\frac{1}{2}\sum_{j=1}^{N}\left(\Big(e^{-\frac{\beta}{4}}|\uparrow\,\uparrow\rangle_{j}-e^{\frac{\beta}{4}}|\downarrow\,\downarrow\rangle_{j}\Big)\bigotimes_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{N}\Big(e^{-\frac{\beta}{4}}|\uparrow\,\uparrow\rangle_{l}+e^{\frac{\beta}{4}}|\downarrow\,\downarrow\rangle_{l}\Big)\right),$
$|\left(S^{z}\right)^{2}\rangle\rangle_{\beta}$ $=\frac{N}{4}\,|{\mathbb{1}}\rangle\rangle_{\beta}+\frac{1}{4}\sum_{i<j}\left(\Big(e^{-\frac{\beta}{4}}|\uparrow\,\uparrow\rangle_{i}-e^{\frac{\beta}{4}}|\downarrow\,\downarrow\rangle_{i}\Big)\otimes\Big(e^{-\frac{\beta}{4}}|\uparrow\,\uparrow\rangle_{j}-e^{\frac{\beta}{4}}|\downarrow\,\downarrow\rangle_{j}\Big)\bigotimes_{\begin{subarray}{c}l=1\\ l\neq i,j\end{subarray}}^{N}\Big(e^{-\frac{\beta}{4}}|\uparrow\,\uparrow\rangle_{l}+e^{\frac{\beta}{4}}|\downarrow\,\downarrow\rangle_{l}\Big)\right).$ (S34)
As noted in the main text, an additional scar of the model (14) can be obtained from the single-layer Hamiltonian $H$. This scar explicitely reads
$|H\rangle\rangle_{\beta}$ $=2\sum_{i<j}w_{ij}\left(\Big(|\uparrow\,\downarrow\rangle_{i}\otimes|\downarrow\,\uparrow\rangle_{j}+|\downarrow\,\uparrow\rangle_{i}\otimes|\uparrow\,\downarrow\rangle_{j}\Big)\bigotimes_{\begin{subarray}{c}l=1\\ l\neq i,j\end{subarray}}^{N}\Big(e^{-\frac{\beta}{4}}|\uparrow\,\uparrow\rangle_{l}+e^{\frac{\beta}{4}}|\downarrow\,\downarrow\rangle_{l}\Big)\right).$ (S35)
Note that these scars have nonzero entanglement for the transverse bipartition, in contrast to the thermofield scar $|{\mathbb{1}}\rangle\rangle_{\beta}$.
Appendix G S.7 A different type of scar states in bilayer systems
Consider a bilayer system with the Hamiltonian
$\mathcal{H}=$ $H_{1}+H_{2}+\sum_{i=1}^{N_{1}}\sum_{j=1}^{N_{2}}g_{ij}\bigl(\sigma^{x}_{i}\tau^{x}_{j}-\sigma^{y}_{i}\tau^{y}_{j}\bigr),$ (S36)
where $H_{1}$ ($H_{2}$) is a Hamiltonian of $N_{1}$ ($N_{2}$) spins $1/2$ constituting the first (second) layer, and the third term in $\mathcal{H}$ is the interlayer interaction. The Hamiltonians $H_{1}$ and $H_{2}$ conserve the $z$-projections $S^{z}_{1}=\sum_{i=1}^{N_{1}}\sigma^{z}_{i}$ and $S^{z}_{2}=\sum_{j=1}^{N_{2}}\sigma^{z}_{j}$ of spins of respective layers. Apart from this, $H_{1}$ and $H_{2}$ are arbitrary (in particular, can contain a different number of spins), and therefore the closed system $\mathcal{H}$ is not covered by the duality discussed in the main text. Still, the states
$|P_{1}\rangle=|\underbrace{\uparrow\uparrow\dots\uparrow}_{1\mathrm{st~layer}}\rangle\otimes|\underbrace{\downarrow\downarrow\dots\downarrow}_{2\mathrm{nd~layer}}\rangle,\qquad|P_{2}\rangle=|\underbrace{\downarrow\downarrow\dots\downarrow}_{1\mathrm{st~layer}}\rangle\otimes|\underbrace{\uparrow\uparrow\dots\uparrow}_{2\mathrm{nd~layer}}\rangle$ (S37)
are eigenstates of $\mathcal{H}$. Indeed, each of this states is a maximal(minimal)-eigenvalue eigenstate of $S^{z}_{1}$ and $S^{z}_{2}$ and hence of $H_{1}$ and $H_{2}$. Further, each of this states is a zero-eigenvalue eigenstate of each term of the form $(\sigma^{x}_{i}\tau^{x}_{j}-\sigma^{y}_{i}\tau^{y}_{j})$, which completes the proof.
In a special case of $\mathcal{H}$ dual to an open system according to eq. (6), the eigenstates $|P_{1}\rangle$ and $|P_{2}\rangle$ are dual to non-Hermitian eigen operators of the Lindbladian
$P_{1}=\prod_{j=1}^{N}\sigma_{j}^{+},\qquad P_{2}=\prod_{j=1}^{N}\sigma_{j}^{-}.$ (S38)